# Strong-field Driven Sub-cycle Band Structure Modulation Measured with Ultrafast Electric Field Observables

### v1

@@ II Results and Discussion
It can be readily seen from Eq. II that the $1\omega_{0}$ peak in Fig. 3 primarily originates from the linear expansion of the net field dependent $\chi^{(3)}$. The second-order expansion contributes not only to the $1\omega_{0}$ peak but also generates a component at $2\omega_{0}$. A similar reasoning extends naturally to the higher-order expansions of $\chi^{(3)}$. Within this framework, the difference in the relative peak strengths at $1\omega_{0}$ and $2\omega_{0}$ can be understood by noting that the third-order susceptibility is predominantly governed by the linear term, i.e., $\chi^{(3)}\approx\chi_{0}^{(3)}+a\mathcal{E}$. Analogous phenomena have previously been observed, where an external field modifies both the linear and nonlinear response functions of a material. A well-known example is the optical Kerr effect, in which the change in the linear refractive index of a medium is proportional to the intensity of the applied field [40, 41, 42]. Similar effects have also been utilized in electric-field-induced second-harmonic generation, where the external field breaks the inversion symmetry of otherwise centrosymmetric systems, making the second-order response function nonzero[43, 44]. However, in this experiment, the observed oscillation cannot be explained by including a fifth-order term due to the phase-matching condition, as explained previously using Eq. II. To further support this interpretation, we isolate the $1\omega_{0}$ peak of the measured chirp in Fig. 3 (b) and inverse Fourier transform to obtain the single frequency filtered oscillation of the temporal chirp, shown in Fig. 4 (a). This is then compared with the $1\omega_{0}$ component of the chirp calculated using Eq. II. Fig. 4 (b) shows the time delay dependent phase of the oscillations in Fig. 4 (a). The model accurately reproduces the $\pi$-phase jump in the chirp oscillation near $\tau=20~\mathrm{fs}$, in excellent agreement with the experimental observation, as shown in Fig. 4 (b). The offset of this phase jump from zero time delay is due to the chirp present in the input fields.
@@ III Conclusion
In summary, we have measured the complete electric field of the degenerate four-wave mixing (DFWM) signal in magnesium oxide using spectral interferometry. By analyzing field observables, such as the time-integrated amplitude and the chirp of the signal, as a function of the sub-cycle time delay between incoming pulses, we identified unique oscillatory features. We attribute these features to the ultrafast modulation of the nonlinear response function of the crystal due to modulation of the band structure. To support this interpretation, we simulated the DFWM experiment using time-dependent perturbation theory and found that oscillatory behavior could only be reproduced when instantaneous ultrafast band structure modulation driven by the strong field is included. Such ultrafast band modulation effects are known to leave signatures in other optical observables, such as high-harmonic spectra [31, 32, 45]. Here, we demonstrate that the influence of field-dressed bands is clearly manifested in the field observables of nonlinear optical signals. Our results indicate that field observables can serve as direct probes of strong-field-driven attosecond electron dynamics in solid-state systems.
Furthermore, we have demonstrated that modeling of our experimental observations requires incorporating strong light–matter interactions into the modification of the third-order nonlinear susceptibility. This finding has significant implications for the manipulation and control of observables governed by the nonlinear response function. For instance, it is well established that signals generated through nonlinear optical processes can serve as sources of non-classical light. Processes such as parametric down-conversion (PDC) and degenerate four-wave mixing (DFWM) are known to produce squeezed light [46, 47, 48, 49, 50]. The degree of squeezing in these processes is directly proportional to the nonlinear coupling coefficients of the respective systems, i.e., $\propto\chi^{(2)}$ for PDC and $\propto\chi^{(3)}$ for DFWM [50]. More recently, it has been demonstrated that squeezing in DFWM signals can also be observed from ultrafast light–matter interactions with few-cycle pulses [51, 52]. With the strong sub-cycle modulation of $\chi^{(3)}$ demonstrated in this work, we believe that it is possible to control the degree of squeezing on attosecond time scales. This could pave the path to achieve extreme levels of squeezing based on field induced nonlinearities that exceed current squeezing levels achieved using periodically-poled nonlinear materials [53] with numerous applications in quantum metrology and quantum information science.
@@ IV.1 Experimental Setup
In our setup of the DFWM experiment, as seen in Fig. 1, three pulses with different propagation vectors interact with the medium in a non-collinear geometry and generate a fourth pulse due to a third-order nonlinear interaction in a BOXCAR configuration. Due to the non-collinear geometry, the signal pulse is spatially separated from the other three beams. The output signal has a momentum of $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. The Probe and Gate beams are centered at 800 nm, with a pulse width of 60 fs and a repetition rate of 1 kHz. The Gate pulse is divided into two phase-locked Gate pulses (Gate 1 and Gate 2) using a mask. The Probe pulse is delayed with respect to the Gate pulses using an optical delay stage, and the Gate polarization can be rotated relative to the Probe polarization. However, in this experiment, the polarizations of all four pulses are fixed along the same direction. The MgO crystal is mounted on a controllable rotation mount to allow for precise control of the crystal angle with respect to the probe polarization. The Probe and the two Gate pulses each have a pulse energy of $\sim$ 500 nJ and form the input pulses in our DFWM measurement. Before the interaction, a Reference pulse is derived from the Probe pulse. After the interaction, the Signal pulse is isolated by an iris, so the residual Probe and Gate beams are blocked. The Signal pulse is passed through a high-contrast polarizer that extracts the polarization component of the Signal that is along the input polarization direction. The output signal pulse was completely measured using TADPOLE [27], where the Reference pulse is combined with the Signal pulse in a spectrometer to obtain the measurement of spectral interference. Frequency-Resolved Optical Gating (FROG) [54] is used to completely characterize the Reference pulse. The measured spectral interference of the DFWM signal and the Reference can be written as:
@@ Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award #\# DE-SC0024234. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility (project m4269). The authors thank Dr. Rudro Biswas for helpful discussions.

### v2

@@ II.2 Strong-field Driven Nonlinear Response Modulation
It can be readily seen from Eq. II.2 that the $1\omega_{0}$ peak in Fig. 3 primarily originates from the linear expansion of the net field-dependent $\chi^{(3)}$. The second-order expansion contributes not only to the $1\omega_{0}$ peak but also generates a component at $2\omega_{0}$. A similar reasoning extends naturally to the higher-order expansions of $\chi^{(3)}$. Within this framework, the difference in the relative peak strengths at $1\omega_{0}$ and $2\omega_{0}$ can be understood by noting that the third-order susceptibility is predominantly governed by the linear term, i.e., $\chi^{(3)}\approx\chi_{0}^{(3)}+a\mathcal{E}$. Analogous phenomena have previously been observed in which an external field modifies both the linear and nonlinear response functions of a material. A well-known example is the optical Kerr effect, in which the change in the linear refractive index of a medium is proportional to the intensity of the applied field [49, 42, 5]. Similar effects have also been utilized in electric-field-induced second-harmonic generation, where the external field breaks the inversion symmetry of otherwise centrosymmetric systems, making the second-order response function nonzero[58, 60]. However, in this experiment, the observed oscillation cannot be explained by including a fifth-order term under the phase-matching condition, as shown previously using Eq. II.2. To further support this interpretation, we isolate the $1\omega_{0}$ peak of the measured chirp in Fig. 3 (b) and inverse Fourier transform to obtain the single frequency filtered oscillation of the temporal chirp, shown in Fig. 4 (a). This is then compared with the $1\omega_{0}$ component of the chirp calculated using Eq. II.2. Fig. 4 (b) shows the time delay-dependent phase of the oscillations in Fig. 4 (a). The model accurately reproduces the $\pi$-phase jump in the chirp oscillation near $\tau=20~\mathrm{fs}$, in excellent agreement with the experimental observation, as shown in Fig. 4 (b). The offset of this phase jump from zero time delay is due to the chirp present in the input fields.
II.3 Sub-cycle control of dephasing
From perturbation theory, the third-order polarization can be written as[4, 31],
$P^{(3)}(t)$ $=\operatorname{Tr}\left[\hat{\mu}\hat{\rho}^{(3)}(t)e^{-t/\tau_{\text{dephase}}}\right]$ (6)
In the density matrix formalism, the diagonal terms correspond to the population of electronic states, and the off-diagonal elements represent coherence between states. As coherences evolve in time phase fluctuations attenuate their temporal evolution, a process called dephasing [31]. To model dephasing of coherences, phenomenological dephasing terms $\gamma_{\text{dephase}}=1/\tau_{\text{dephase}}$ are added. This means that the output-field pulse width depends on the dephasing of the coherence created by the probe pulse in the final step of the nonlinear interaction. To quantify third-order coherence dephasing, we fit the measured output temporal field with an Exponentially Modified Gaussian (EMG) function. We fix the Gaussian width ($\sigma$) to 21.2 fs (corresponding to the input probe pulse FWHM of 50 fs) and extract the dephasing time $\tau$ of the DFWM signal field as shown in Fig. 5 (b) (see supplementary materials for details on fitting). The extracted dephasing times $\tau_{\text{dephase}}$ from the EMG fit are plotted as a function of time delay in Fig. 5 (c). The data indicate that band-structure modification resulting from strong light-matter interaction modulates the dephasing times of nonlinear interactions on sub-cycle time scales. Multiple factors could influence dephasing times in solids, including doping, electron-electron interactions, and electron-phonon interactions. In this study, the measured dephasing time likely reflects electron-electron interactions that affect real-space coherence of electron-hole pairs across lattice sites, as in the case of HHG [7]. Furthermore, band-structure modulation induced by strong light-matter interaction also alters electron-electron interactions, thereby modulating dephasing times.
@@ III Conclusion
In summary, we have measured the complete electric field of the degenerate four-wave mixing (DFWM) signal in magnesium oxide using spectral interferometry. By analyzing field observables, such as the time-integrated amplitude and the signal field chirp, as functions of the sub-cycle time delay between incoming pulses, we identified unique oscillatory features. We attribute these features to the ultrafast modulation of the nonlinear response function of the crystal due to modulation of the band structure. To support this interpretation, we simulated the DFWM experiment using time-dependent perturbation theory and found that oscillatory behavior could be reproduced only when instantaneous ultrafast band-structure modulation driven by the strong field is included. Such ultrafast band modulation effects are known to leave signatures in other optical observables, such as high-harmonic spectra [52, 53, 25]. Here, we demonstrate that the influence of field-dressed bands is clearly manifested in the field observables of nonlinear optical signals. Our results indicate that field observables can serve as direct probes of strong-field-driven attosecond electron dynamics in solid-state systems. We have demonstrated that modeling of our experimental observations requires incorporating strong light–matter interactions into the non-perturbative modification of the third-order nonlinear susceptibility. Furthermore, by fitting the real-time electric field envelope of the signal pulses, we have extracted dephasing times corresponding to this non-resonant nonlinear interaction. This represents a direct measurement of dephasing in a nonlinear process via electric field measurement opening doors to enhacing our understanding of dephasing mechanisms. The modulation of the band structure is seen to modulate the dephasing time, resulting in sub-cycle control of dephasing.
Our findings have significant implications for the manipulation and control of observables governed by the nonlinear response function. For instance, it is well established that signals generated through nonlinear optical processes can serve as sources of non-classical light. Processes such as parametric down-conversion (PDC) and degenerate four-wave mixing (DFWM) are known to produce squeezed light [3, 6, 1, 35, 12]. The degree of squeezing in these processes is directly proportional to the nonlinear coupling coefficients of the respective systems, i.e., $\propto\chi^{(2)}$ for PDC and $\propto\chi^{(3)}$ for DFWM [12]. It has been demonstrated that squeezing in DFWM signals can also be observed from ultrafast light–matter interactions [40, 45, 62]. With the strong sub-cycle modulation of $\chi^{(3)}$ demonstrated in this work, it is now possible to control the degree of squeezing on attosecond time scales [62]. This could pave the path to achieve extreme levels of squeezing based on field induced nonlinearities that exceed current squeezing levels achieved using periodically-poled nonlinear materials [54] with numerous applications in quantum metrology, quantum spectroscopy, and quantum information science.
@@ IV.1 Experimental Setup
In our setup of the DFWM experiment, as seen in Fig. 1, three pulses with different propagation vectors interact with the medium in a non-collinear geometry and generate a fourth pulse due to a third-order nonlinear interaction in a BOXCAR configuration. Due to the non-collinear geometry, the signal pulse is spatially separated from the other three beams. The output signal has a momentum of $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. The Probe and Gate beams are centered at 800 nm, with a pulse width of 50 fs and a repetition rate of 1 kHz. The Gate pulse is divided into two phase-locked Gate pulses (Gate 1 and Gate 2) using a mask. The Probe pulse is delayed with respect to the Gate pulses using an optical delay stage, and the Gate polarization can be rotated relative to the Probe polarization. However, in this experiment, the polarizations of all four pulses are fixed along the same direction. The MgO crystal is mounted on a controllable rotation mount to allow precise control of the crystal angle relative to the probe polarization. The Probe and the two Gate pulses each have a pulse energy of $\sim$ 500 nJ and form the input pulses in our DFWM measurement. Before the interaction, a Reference pulse is derived from the Probe pulse. After the interaction, the Signal pulse is isolated by an iris, so the residual Probe and Gate beams are blocked. The Signal pulse passes through a high-contrast polarizer, which extracts the polarization component of the signal along the input polarization direction. The output signal pulse was completely measured using TADPOLE [13], where the Reference pulse is combined with the Signal pulse in a spectrometer to obtain the measurement of spectral interference. Frequency-Resolved Optical Gating (FROG) [51] is used to completely characterize the Reference pulse. The measured spectral interference of the DFWM signal and the Reference can be written as:
@@ Acknowledgments
This work was primarily supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award #\# DE-SC0024234. NS acknowledges support from the National Science Foundation under award #\# 2208061, which supported YZ. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility (project m4269). The authors thank Dr. Rudro Biswas for helpful discussions.